how to calculate this problem: p=10 Tca=70+30Q+2Q'2 Tcb=400-50Q+5Q'2 find 1 profit 2 quantity
1
Expert's answer
2019-08-27T09:40:18-0400

For competitive market: P= MC

MC = (TC)' = (70+30Q+2Q'2+400-50Q+5Q'2)' = (470 -20Q +7Q'2)' = -20+14Q

10 = -20+14Q

14Q = 30

Q = 2 - quantity

2*10 = 20 - profit